OpenStudy (anonymous):

The formula for converting the temperature in Fahrenheit (F) to Celsius (C) is C = 5/9 (F - 32) . Find F when C = –40 F = –40 F = –14.4 F = –54 F = 9.7

OpenStudy (anonymous):

[-40/(-5/9)]+32=F can u solve for that??

OpenStudy (anonymous):

its 5/9,not negative i mean..

OpenStudy (anonymous):

it took me right back to -40 and if i subtract i get -104

OpenStudy (anonymous):

no the answer is -40

OpenStudy (anonymous):

ok i got the answer its 9.7

OpenStudy (anonymous):

-40/(5/9)+32=F -72+32=F -40=F or F=-40

OpenStudy (anonymous):

how did you get 9.7?? can u show ur solution??